OpenStudy (anonymous):

For what value of the variable is the reciprocal undefined? 2x/15

OpenStudy (anonymous):

The reciprocal simply put as flipping the numerator and the denominator. Example: the reciprocal of 3/5 is 5/3.

OpenStudy (anonymous):

So for which values of x would 15/2x be undefined?

OpenStudy (anonymous):

The reciprocal of 2x/15 would be 15/2x, and if x = 0, 2x = 2(0) = 0, making 15/2x = 15/0 which results in an answer that is undefined.